OpenStudy (anonymous):

Help please? Please help me with widest to narrowest: y=x^2, y=5x^2, y=3x^2.

OpenStudy (anonymous):

Why not graph all of them and see for yourself wich is the widest, etc..

OpenStudy (anonymous):

Wait! How do I graph them?

OpenStudy (anonymous):

Do you have agraphing calculator?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Did you learn how to graph y = x^2 in class?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

How did your teacher assign such a question if you didnt learn how to graph y = x^2 in class beats me.

OpenStudy (anonymous):

I don't know either. This is very new to me.

OpenStudy (anonymous):

But in any event, if you take a function, say y = x^2 and multiply it by a number greater than 1, say 5, you get 5x^2 which will be narrower than y = x^2. So 5x^2 is the narrowest of the three; 3x^2 will be narrower than x^2. So from the three, 5x^2 is the narrowest, then 3x^2, then x^2.

OpenStudy (anonymous):

y = x^2 is a U-shaped parabola

OpenStudy (anonymous):

so 5x^2 will also be U-shaped nut narrower than x^2

OpenStudy (anonymous):

but*

OpenStudy (anonymous):

Ah, thank you so much! I think I understand it now.

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

Thank you for the help^^

OpenStudy (anonymous):

welcome.